# Multivalued (Wess-Zumino-Novikov) Functional in Fluid Mechanics

### v1

@@ Abstract
A homentropic flow occurs when the density ratio $S$ is uniform and constant. In this situation, and also in barotropic flows the vorticity tensor is degenerate ${\rm\ det}\,\Omega_{\mu\nu}=0$, having rank 2. This prevents the construction of the WZN-term since the phase space of a homentropic flow is not symplectic. The homentropic flow conserves helicity $\Sigma^{\mu}$, and independently conserves the particle number $n^{\mu}$ providing no exchange with the reservoir. In this case, the equations motion are no different from that of the perfect fluid. This can be observed from the relation $\Sigma^{\lambda}\partial_{\lambda}S=\tfrac{1}{2}\left({d\bar{n}}/{dn}\right)_{\varepsilon}\,\partial_{\mu}\Sigma^{\mu}$ which follows from (45) [the derivative is taken at a fixed energy]. When $S$ is constant, both $\partial_{\mu}\Sigma^{\mu}$ and the WZN term vanish.
The work was supported by the NSF under Grant NSF DMR-1949963. He acknowledges insightful discussions with V. P. Nair, G. Volovik, M. Stephanov, B. Shapiro, L. Friedlander, and is especially grateful to A. G. Abanov and A. Cappelli for collaboration on this subject.

### v2

@@ Abstract
A homentropic flow occurs when the density ratio $S$ is uniform and constant. In this situation, and also in barotropic flows the vorticity tensor is degenerate ${\rm\ det}\,\Omega_{\mu\nu}=0$, having rank 2. This prevents the construction of the WZN-term since the phase space of a homentropic flow is not symplectic. The homentropic flow conserves helicity $\Sigma^{\mu}$ and independently conserves the particle number $n^{\mu}$ providing no exchange with the reservoir. In this case, the equations of motion are no different from that of the perfect fluid. This can be observed from the relation for the potential vorticity $nwe:=\Sigma^{\lambda}\partial_{\lambda}S=\tfrac{1}{2}\left({d\bar{n}}/{dn}\right)_{\varepsilon}\,\partial_{\mu}\Sigma^{\mu}\,,$ which follows from (45) [here the derivative is taken at a fixed energy]. When $S$ is constant, both $\partial_{\mu}\Sigma^{\mu}$ and the WZN term vanish.
Acknowledgements.
The work was supported by the NSF under Grant NSF DMR-1949963. The author gratefully acknowledges discussions V. P. Nair, G. Volovik, M. Stephanov, B. Shapiro, L. Friedlander. Special appreciation is extended to A. G. Abanov and A. Cappelli for their collaboration on this subject.